OpenStudy (anonymous):

Please do you know: The length of a rectangle is 3yd more than twice its width and the area of the rectangle is 77y2 find the dimension of the rectangle

OpenStudy (anonymous):

let width b=x length l=2x+3 area=l*b x(2x+3)=77 solve and get the solution.

OpenStudy (anonymous):

What is the length and width for the quadratic equations

OpenStudy (anonymous):

\[2 x ^{2}+3x=77, or2x ^{2}+3x-77=0\] 2*-77=-154 14-11=3 14*-11=-154 \[x ^{2}+\left( 14-11 \right)x-77=0\] \[x ^{2}+14x-11x-77=0\] make factors and find x i.,b and l

OpenStudy (anonymous):

\[i wrote wrong please 2 with x ^{2}\] \[2x ^{2}+14x-11x-77=0\] \[2x \left( x+7 \right)-11\left( x+7 \right)=0\] \[\left( x+7 \right)\left( 2x-11 \right)=0,either x=-7 or2x=11,x=\frac{ 11 }{2 }\] width =11/2 yards length=2*11/2+3=14yards

OpenStudy (anonymous):

because width or length can not be negative

OpenStudy (anonymous):

Ok. Thanks

OpenStudy (anonymous):

yw